# Beehive: A Flexible Network Stack for Direct-Attached Accelerators

### v3

@@ Abstract.
Accelerators have become increasingly popular in datacenters due to their cost, performance, and energy benefits. Direct-attached accelerators, where the network stack is implemented in hardware and network traffic bypasses the main CPU, can further enhance these benefits. However, modern datacenter software network stacks are complex, with interleaved protocol layers, network management functions, as well as virtualization support. They also need to flexibly interpose new layers to support new use cases. By contrast, most hardware network stacks only support basic protocol compatibility and are often difficult to extend due to using fixed processing pipelines.
This paper proposes Beehive, a new, open-source hardware network stack for direct-attached FPGA accelerators designed to enable flexible and adaptive construction of complex protocol functionality. Our approach is based on a network-on-chip (NoC) substrate, automated tooling for the independent scale-up of protocol elements, compile-time deadlock analysis, and a flexible diagnostics and control plane. Our implementation interoperates with standard Linux TCP and UDP clients, allowing existing RPC clients to interface with the accelerator. We use three applications to illustrate the advantages of our approach: a throughput-oriented erasure coding application, an accelerator for distributed consensus operations that reduces the latency and energy cost of linearizability, and TCP live migration support for dynamic server consolidation.
1. Introduction
Figure 1. Box (a) represents the standard CPU server node. Box (b) represents Beehive’s targeted, direct-attached setup where Beehive provides the network stack. Box (c) represents the traditional method of attaching an accelerator to the network using a CPU network stack. Beehive enables direct-attached accelerators to interoperate with other types of nodes on the network.
@@ 1. Introduction
Hardware accelerators are becoming increasingly common in datacenters to reduce cost, improve performance, and reduce energy consumption relative to server CPUs. Typically, accelerators are hosted over the PCIe I/O bus, with the server CPU mediating all communication with the accelerator, illustrated in Box (c) in Figure 1. An emerging alternative model directly attaches the accelerator to the network, with its own network functionality implemented in hardware, illustrated in Box (b) in Figure 1. By bypassing the CPU, such a specialized hardware network stack reduces latency, performance variability, and overhead, freeing up the CPU for other purposes. Another advantage of this approach is that it simplifies scaling out accelerators to meet workload needs, independently of server hardware.
@@
A key barrier to fully realizing the benefits of direct-attached hardware acceleration is the growing disparity between datacenter networking requirements and hardware network stack implementations. Modern datacenter networks have intricate host network stacks that are constantly evolving to meet customer and operator use cases (google_snap, ). Beyond core protocols, such as TCP/IP, modern applications require higher-level host network functionality like remote procedure call (RPC) processing, quality-of-service (QoS) management (aequitas, ; breakwater, ), encryption (ktls, ; mtls, ), and load balancing (istio_l7_lb, ; cilium_l7_lb, ). Deployment flexibility necessitates management features like virtual networking (openvswitch_network_virt_nsdi, ; microsoft_vfp_nsdi, ; google_andromeda_nsdi, ), access control lists (hyperv_acl, ), congestion control (swift, ; hpcc, ), traffic prioritization (homa, ; bfc, ), and load balancing (ananta_sigcomm, ; maglev_nsdi, ; facebook_katran, ; cloudflare_unimog, ). Deployment maintainability requires dynamic support for network monitoring (pigasus_osdi, ; ballard2010extensible, ), reconfiguration (rmt_sigcomm, ; kim2013improving, ), and debugging (tammana2016simplifying, ).
@@
An example of a highly-flexible software network stack is Google’s Snap networking system (google_snap, ). It is designed around composable message-passing engines, with modules for load balancing, network virtualization, network management, and custom transport protocols. New modules can be easily inserted anywhere in the stack, without re-engineering the rest of the stack. By contrast, existing hardware network stacks are specialized and rigid since they are typically designed to support a single application with few protocol layers. Although some work focuses on flexible packet-level processing in hardware (clicknp_sigcomm, ; panic_osdi, ), they operate at the packet handling layer with no support for higher-level transport and application protocols. Other work focuses on hardware transport offloads, but these systems lack a range of essential network functions (limago, ; chelsio_t6, ; tonic, ).
@@
This paper explores the design of an FPGA network stack for direct-attached accelerators that can better support the extensibility and complexity needed in a production environment. Flexibility is needed at multiple layers of the network stack: in packet processing (layer 3), transport and congestion control (layer 4), and the application layer (layer 7). Adding new functionality, scaling out a protocol layer to meet workload demand, or changing internal load balancing should be simple, as it is in software, without disrupting or re-engineering other layers. Our goal is to make a hardware network stack that is flexible and scalable while preserving the cost, performance, and energy benefits of direct-attached accelerators.
@@
We propose and implement Beehive, an open-source hardware network stack architected as a collection of protocol layers and network functions that communicate via message-passing over a scalable network-on-chip (NoC). We provide automated tooling for managing scale out, load balancing, and compile-time deadlock analysis. To make our design concrete, we implement Ethernet, IP, UDP, TCP, network address translation (NAT), IP-in-IP encapsulation, and additional support for control and debugging of network functions. Our implementation interoperates with Linux TCP and UDP clients, allowing unmodified remote procedure call (RPC) clients to use our accelerator.
For our evaluation, we implement Beehive on an FPGA and show that it offers up to 31x higher per-core throughput than state-of-the-art CPU kernel-bypass networking stacks on small messages. We also demonstrate how Beehive can improve performance and energy consumption in three important use cases compared to CPU-only implementations.
@@
First, modern data center storage systems typically use erasure coding to achieve better storage efficiency than replication with comparable fault tolerance. We implement an erasure coding accelerator in Beehive and show that, compared to a CPU-only version, the accelerator scales out to 62 Gbps using  20$\times$ less energy.
Second, many production distributed systems use consensus algorithms for data consistency even in case of failures. However, deployed systems often skip consensus for some reads in exchange for better performance. We show that accelerating a key piece of the consensus algorithm in hardware can reduce end-to-end median operation latency by 1.13$\times$, with 1.14$\times$  better per-core throughput and  2$\times$ less energy than the CPU-only version. Because distributed consensus is dominated by small messages, it requires independent scale out and load balancing of different protocol layers, implemented without changing the protocol logic.
Finally, we demonstrate the advantages of Beehive’s message-passing model to add support for TCP live migration by adding network address translation and management capabilities without changes to any other protocol layers. This design allows accelerator clients to be transparently migrated without resetting their TCP connection, with less than a millisecond of added end-to-end operation delay.
@@
  • An open-source FPGA implementation of Beehive that includes tools and reusable components to build network stacks for accelerators that use different transport protocols, network virtualization, and L7 functionality.
  • A demonstration of Beehive’s ability to support scalability, flexibility, low latency, high throughput, and energy efficiency by integrating and evaluating an erasure coding accelerator, a consensus accelerator, and TCP migration functionality.
Figure 2. A high-level diagram of the type of network stack Beehive targets. Along with multiple transport protocols, this stack has IP-in-IP and VXLAN for network virtualization and a component for an L4 load balancer. The control plane potentially needs access to all components.
2. Beehive Design Goals
Our overarching goal for Beehive is to build an open-source FPGA hardware design to support emerging applications for direct network-attached accelerators in a production environment. Figure 2 shows a high-level diagram of the type of network stack architecture we want to be able to support. Applications may only use some subset of these protocols and network functions. This requires:
@@ 2. Beehive Design Goals
Modularity. However, network stacks are not fixed. Requirements are constantly changing with new custom protocols (e.g. Google’s Pony Express (google_snap, ) or 1RMA (1rma_sigcomm, )) and network functions. In order to facilitate rapid development and customization of the network stack, our framework must be modular, so we can compose or integrate new components with minimal to no modifications to existing components.
@@
Management flexibility. Components in a network stack need to be able to interact beyond just passing packet data. For example, components need to be able to expose interfaces to the control plane for telemetry and debugging (envoy, ). The control plane may also need to update state used by a protocol or network function, such as configuring the load balancer used to parcel work across application accelerator instances. Such configurability should be possible even in large designs without extensive manual optimization.
3. Design
Figure 3. Architecture of a tile.
3.1. Beehive’s Architecture
The basic component in Beehive is the tile, shown in Figure 3. Each tile has a network-on-chip (NoC) router, some logic that handles NoC message construction and deconstruction, and some processing logic, such as a protocol layer, network function, or application. Tile routers are connected together to form the NoC topology. We do not require a particular topology, although our prototype uses a 2D mesh. We assume the NoC is reliable, point-to-point ordered, and uses deterministic, deadlock-free routing.
@@ 3.1. Beehive’s Architecture
Each tile hop is responsible for determining the next tile that a message should be sent to. This design is in contrast to earlier work which assumes that routes can be fully determined on packet arrival (panic_osdi, ). We discuss this decision in more detail in Section 3.4. This component may vary in complexity from a static CAM to more complex logic, such as content-based routing. The set of possible message chains is known ahead of time for deadlock analysis ,described in Section 3.5.
@@
Figure 4. The flow through which a packet is processed or constructed in Beehive.
3.2. Processing a Packet
Figure 4 shows an example of a basic UDP stack in Beehive, with a UDP packet moving through the receive and send paths.
@@ 3.2. Processing a Packet
On the receive side, an Ethernet frame enters the Ethernet tile, which has ports for the I/O from the transceivers in addition to the ports connecting to other tiles. The processing logic within the tile parses and removes the Ethernet header, realigning the data. This is then turned into a NoC message consisting of a header flit, a metadata flit with the parsed Ethernet header, and some number of data flits containing the remaining packet data. The routing component in the Ethernet tile uses the type field in the Ethernet header to determine that the message should be passed to the IP tile. The IP tile similarly parses the IP header, validates the header’s checksum, and then creates a NoC message to be sent to the UDP layer based on the protocol field. Finally, the UDP tile parses the UDP header, validates the packet’s checksum, and generates a NoC message to be sent to the application based on the port in the UDP header. The transmit path runs similarly, except instead of parsing headers from the data flits, headers are added into the data flits by each protocol tile. After the Ethernet tile adds on the Ethernet header, it is sent out the ports for I/O with the transceivers. This incremental composability is good for our goal of modularity as it makes it easier to insert new functionality between stages.
@@
While there is only one possible destination for the tiles in this design, there can potentially be multiple endpoints, such as other protocols (e.g. TCP connected to IP), network services (e.g. network virtualization), or replicated tiles for higher bandwidth. With replicated tiles, there are multiple ways to decide on which tile should receive an incoming packet. The simplest method is to distribute packets between them in a round-robin fashion. However, more complex scheduling may be necessary if a tile holds state for particular flow. In this case, it is important that packets from the same flow always go to the same tile. This distribution can either be integrated within a tile or placed in a dedicated tile. We discuss examples about how we distribute packets to duplicated tiles in Section 5.
3.3. Message-Passing Interconnect
@@
Being able to compose elements is essential for facilitating customization. We opt for a message passing model. This is beneficial for modularity, because defining a message-passing format allows us to standardize the physical interconnection between components, a recognized benefit in SoC design (route_packets_dac, ), and makes it easier to chain offloads together. ClickNP (clicknp_sigcomm, ) and PANIC (panic_osdi, ), two modular packet processing frameworks, have also used a message-based approach. The message passing can be done over dedicated connections, which is the approach used by ClickNP, or a NoC which is used by PANIC.
@@
We prefer a NoC interconnect for two main reasons related to our goal of scalability. First, we can take advantage of the multiplexing that is provided by the NoC routers. Certain tiles may interact with a large number of other tiles, e.g. if we instantiate multiple copies of the same component. Direct connections can lead to large multiplexers and wires with significant fan-out. Although we could create specialized pipelined multiplexers and arbiters, these essentially look like NoC routers.
@@
Second, we would like the interconnect wiring to remain stable whenever possible. In the ClickNP model, top-level wires are determined by the computational graph. If we wish to form a chain that links together two components that did not communicate before, we must add new interconnect wires, which are typically the longest wires. A NoC allows us to reuse physical wiring to chain any elements that exist in the design, although we must be careful with deadlock.
3.4. Tile Chain Routing

### v5

@@ Abstract
Direct-attached accelerators, where application accelerators are directly connected to the datacenter network via a hardware network stack, offer substantial benefits in terms of reduced latency, CPU overhead, and energy use. However, a key challenge is that modern datacenter network stacks are complex, with interleaved protocol layers, network management functions, and virtualization support. To operators, network feature agility, diagnostics, and manageability are often considered just as important as raw performance. By contrast, existing hardware network stacks only support basic protocols and are often difficult to extend since they use fixed processing pipelines.
We propose Beehive, a new, open-source FPGA network stack for direct-attached accelerators designed to enable flexible and adaptive construction of complex network functionality in hardware. Application and network protocol elements are modularized as tiles over a network-on-chip substrate. Elements can be added or scaled up/down to match workload characteristics with minimal effort or changes to other elements. Flexible diagnostics and control are integral, with tooling to ensure deadlock safety. Our implementation interoperates with standard Linux TCP and UDP clients, with a 4x improvement in end-to-end RPC tail latency for Linux UDP clients versus a CPU-attached accelerator. Beehive is available at https://github.com/beehive-fpga/beehive
Index Terms:
hardware acceleration, networking, network stack, FPGA
I Introduction
@@
Hardware accelerators are becoming increasingly common in datacenters to reduce cost, improve performance, and reduce energy consumption relative to server CPUs. Typically, accelerators are hosted over the PCIe I/O bus, with the server CPU mediating all communication with the accelerator, illustrated in Figure 1(c). An alternative model directly attaches the accelerator to the network, with its own network functionality implemented in hardware, illustrated in Figure 1(b). Bypassing the CPU potentially reduces end-to-end latency, latency variability, and overhead, freeing up the CPU for other purposes.
@@
A barrier to any hardware network implementation is the difficulty of meeting the full set of datacenter network operational requirements [53, 6]. Network manageability, diagnostic visibility, and interoperability are often non-negotiable requirements, made more complex by the rapid evolution in host network stacks to meet application and operational needs. Beyond core protocols, such as TCP/IP, modern applications require higher-level functionality like remote procedure call (RPC) processing, quality-of-service (QoS) management [80, 17], encryption [50, 19], application-specific load balancing [36, 18], and information flow control [28]. Deployment flexibility necessitates management features like virtual networking [44, 29, 23], access control lists [54], congestion control [45, 48], traffic prioritization [57, 33], and load balancing [61, 26, 70, 75]. Deployment maintainability requires dynamic support for network monitoring [81, 8], reconfiguration [11, 43], and debugging [72].
Figure 1: (a) represents a standard CPU server node; (b) a direct-attached accelerator using the Beehive network stack; (c) an accelerator using a CPU network stack.
@@
An example of a highly-flexible software network stack is Google’s Snap networking system [53]. It is designed around composable message-passing engines, with modules for load balancing, network virtualization, network management, and custom transport protocols. New modules can be easily inserted anywhere in the stack, without re-engineering the rest of the stack. Our question is whether we can do something similar in hardware. Existing hardware network stacks are typically designed to support only a single application with minimal protocol complexity. Although some recent work has focused on flexible packet-level processing in hardware [47, 49], our aim is to support flexibility across the entire network stack, including transport and application protocols. Other work has looked at hardware offload of transport protocols, but these systems lack a range of essential network functions [66, 16, 4], or in the case of RDMA, require extensive engineering to make work in practice [6, 65].
@@
This paper explores the design of an FPGA network stack that can realize the benefits of direct-attached accelerators while supporting the extensibility, incremental scalability, and manageability needed for production use. Flexibility is needed at multiple points in the network stack: in packet processing (layer 3), transport and congestion control (layer 4), the application layer (layer 7), and in control/diagnostics operating alongside, and using, the data plane. Adding new functionality, differentially scaling protocol elements to meet application throughput needs, or inserting a new load balancing policy should be simple, as it is in software, without the need to disrupt or re-engineer other layers.
@@
We propose and implement Beehive, an open-source hardware network stack architected as a collection of protocol functions that communicate via message-passing over a scalable network-on-chip (NoC). We provide automated tooling for managing differential scaling and load balancing of protocol elements, a control plane for diagnostics monitoring, and compile-time deadlock analysis. To make our design concrete, we implement Ethernet, IP, UDP, TCP, network address translation (NAT), IP-in-IP encapsulation, and additional support for control and debugging of network functions. Our implementation interoperates with Linux TCP and UDP clients, allowing unmodified remote procedure call (RPC) clients to use our accelerator.
For our evaluation, we implement Beehive and evaluate it on FPGA. We show that it offers a 4$\times$/1.5$\times$ improvement in end-to-end client RPC tail latency over Linux/user-level TCP relative to mediating accelerator traffic through the server CPU, and up to 31$\times$ higher per-core throughput than a state-of-the-art CPU kernel-bypass stack on small messages.
@@
We implement two example applications using Beehive: erasure coding as a bandwidth-oriented application and distributed consensus as a latency-sensitive application. First, modern datacenter storage systems often use erasure coding for better storage efficiency than replication with comparable fault tolerance. We implement an erasure coding accelerator in Beehive and show that, compared to a CPU-only version, the accelerator scales out to 62 Gbps using 20$\times$ less energy. Second, we show that accelerating a key piece of distributed consensus in hardware can reduce end-to-end median operation latency by 1.13$\times$, with 1.14$\times$  better per-core throughput and  2$\times$ less energy than the CPU-only version.
Figure 2: A high-level diagram of the type of network stack Beehive targets. Along with multiple transport protocols, this stack has IP-in-IP and VXLAN for network virtualization and a component for an L4 load balancer. The downward arrows represent control-plane communication, which potentially needs access any module internal to the network stack.
@@
  • An open-source FPGA implementation of Beehive that includes tools and reusable components to build network stacks for accelerators that use different transport protocols, network virtualization, and layer 7 functionality.
  • A demonstration of Beehive’s ability to support scalability, flexibility, low latency, high throughput, and energy efficiency by integrating and evaluating an erasure coding accelerator and a consensus accelerator.
II Motivation
We motivate direct-attached accelerators by investigating their latency benefits over CPU-attached accelerators. Prior work has shown benefits over the Linux network stack [13, 74]. However, cutting-edge systems aiming for the lowest possible latency typically use a DPDK network stack, which can achieve single digit microsecond latencies [79, 39, 73].
Our experiment compares the direct-attached configuration in Figure 1(b) and the software-hosted configuration in Figure 1(c). We evaluate the performance of UDP echo, where the client sends a UDP packet to a server and waits for the response packet before sending another. We use Linux and F-Stack [73], a DPDK network stack, as software network stacks. We run 1,000,000 requests and measure the round-trip time (RTT) for each request.
For the direct-attached configuration, we use Beehive implementing a UDP echo server. We try both Linux and F-Stack as the clients. For the software-hosted configuration, we use either the Linux network stack or F-Stack as the software network stack and Ensō [67] as the FPGA accelerator. Ensō is an FPGA-based NIC designed for efficient NIC-CPU communication over PCIe. Internally, we tie Ensō’s network output to its input, so it operates as a loopback. For software-hosted configurations, the client and server machines run the same software stack (Linux or F-Stack).
TABLE I: Comparison of median and p99 round-trip time of a UDP echo across different configurations. Client machines use software networking. Beehive represents the configuration in Figure 1(b); Linux and DPDK to Accel. represent Figure 1(c).
Client Linux Client DPDK Client
Server Beehive
Linux
to Accel.
Beehive
DPDK
to Accel.
Median Latency ($\mu$s) 11.6 17.6 4.08 6.22
p99 Latency ($\mu$s) 15.3 61.2 4.43 6.79
We report median and 99th percentile (p99) RTTs in Table I. As expected, trampolining every RPC through the CPU on the way to the FPGA is both slower, and more variable, than when the FPGA is directly attached to the network using Beehive. When the network stack is provided by Linux, message latency can be affected by CPU scheduling contention, so that Beehive has 4$\times$ better p99 tail latency than redirection through the CPU on this benchmark, and 1.5$\times$ better median latency. When the network stack is at user level on both the client and server, scheduling variance is reduced as the server CPU busy-waits for incoming requests, at the cost of higher CPU overhead. However, the relative benefit of Beehive is similar, with 1.5$\times$ better median and p99 tail latency relative to redirection through the CPU.
This shows that even with a DPDK stack, direct-attached accelerators can still provide a latency improvement, and the relative improvement is larger for tail latency compared to the Linux network stack. With this in mind, direct-attached accelerators are an appealing option. Realizing this benefit requires a hardware network stack that can be flexibly reconfigured to meet the needs of datacenter network management.
III Design Goals
Our overarching goal for Beehive is to build an open-source FPGA hardware design to support emerging applications for direct network-attached accelerators in a production environment. Figure 2 shows a high-level diagram of the type of network stack architecture we want to be able to support. Applications may only use some subset of these protocols and network functions. We now discuss our specific design goals.
III-A Beehive Goals
@@
Modularity. However, network stacks are not fixed. Requirements are constantly changing with new custom protocols (e.g. Google’s Pony Express [53] or 1RMA [2]) and network functions. In order to facilitate rapid development and customization of the network stack, our framework must be modular, so we can compose or integrate new components with minimal to no modifications to existing components.
@@
Management flexibility. Components in a network stack need to be able to interact beyond just passing packet data. For example, components need to be able to expose interfaces to the control plane for telemetry and debugging [27]. The control plane may also need to update state used by a protocol or network function, such as configuring the load balancer used to parcel work across application accelerator instances. Such configurability should be possible even in large designs without extensive manual optimization.
III-B Comparing versus related work
As shown in Table II, other related work does not meet all these goals. In terms of complexity, the Limago, a TCP engine written in Vitis HLS, is the closest to Beehive. However, it is not designed to allow for addition or replication of components within the stack, so it is limited in scalability and modularity. We discuss FPGA utilization comparisons further in Section VII-G. Unfortunately, we were unable to run Limago on FPGA using their code [32] to evaluate its performance, because the QSFPs did not come up on the FPGA board.
PANIC and ClickNP are the most similar architecturally to Beehive as they are both based on message-passing over an interconnect, leading to similar performance and modularity benefits as Beehive. Their implementations do not provide standard protocol support directly, but they could be extended to support the logic needed for these protocols. Additionally, their interconnects can limit their scalability. While working on the experiment in Section VII-C, we found PANIC’s crossbar was unable to support more than 8 endpoints, 4 of which are always used by its infrastructure. In ClickNP, components are directly connected using FIFOs, potentially causing fan-out issues when duplicating components. Because ClickNP is not open-source, we were unable to compare to it directly.
Std.
Protocols
Modular Scalable Performant
Mgmt.
Features
Open
Source
Limago [66]
PANIC [49]
ClickNP [47]
LTL [13]
Beehive
TABLE II: Beehive and prior work versus the goals in Section III-A. The stars indicate partially meeting the goal.
IV Design
Figure 3: Architecture of a tile.
IV-A Beehive’s architecture
The basic component in Beehive is the tile, shown in Figure 3. Each tile has a network-on-chip (NoC) router, some logic that handles NoC message construction and deconstruction, and some processing logic, such as a protocol layer, network function, or application. Tile routers are connected together to form the NoC topology. We do not require a particular topology, although our prototype uses a 2D mesh. We require that the NoC is reliable, point-to-point ordered, and uses deterministic, deadlock-free routing.
@@ IV-A Beehive’s architecture
Each tile hop is responsible for determining the next tile that a message should be sent to. This design is in contrast to earlier work which assumes that routes can be fully determined on packet arrival [49]. We discuss this decision in more detail in Section IV-D. This component may vary in complexity from a static CAM to more complex logic, such as content-based routing. The set of possible message chains is known ahead of time for deadlock analysis, described in Section IV-E.
@@
Figure 4: The flow through which a packet is processed or constructed in Beehive.
IV-B Processing a packet
@@
Figure 4 shows an example of a basic UDP stack in Beehive, with a UDP packet moving through the receive and send paths. On the receive side, an Ethernet frame enters the Ethernet tile, which has ports for the I/O from the transceivers in addition to the ports connecting to other tiles. The processing logic within the tile parses and removes the Ethernet header, realigning the data. This is then turned into a NoC message consisting of a header flit, a metadata flit with the parsed Ethernet header, and some number of data flits containing the remaining packet data. The routing component in the Ethernet tile uses the type field in the Ethernet header to determine that the message should be passed to the IP tile. The IP tile similarly parses the IP header, validates the header’s checksum, and then creates a NoC message to be sent to the UDP layer. Finally, the UDP tile parses the UDP header, validates the packet’s checksum, and generates a NoC message to be sent to the application based on the port in the UDP header. The transmit path runs similarly, except instead of parsing headers from the data flits, headers are added by each protocol tile. After the Ethernet tile adds on the Ethernet header, it is sent out the ports for I/O with the transceivers. This incremental composability is good for our goal of modularity as it makes it easier to insert new functionality between stages.
@@
While there is only one possible destination for the tiles in this design, there can potentially be multiple endpoints, such as other protocols (e.g. TCP connected to IP), network services (e.g. network virtualization), or replicated tiles for higher bandwidth. With replicated tiles, there are multiple ways to decide on which tile should receive an incoming packet. The simplest method is to distribute packets between them in a round-robin fashion. However, more complex scheduling may be necessary if a tile holds state for particular flow. In this case, it is important that packets from the same flow always go to the same tile. This distribution can either be integrated within a tile or placed in a dedicated tile. We discuss how we distribute packets to duplicated tiles in Section VI.
IV-C Message-passing interconnect
@@
Being able to compose elements is essential for facilitating customization. We opt for a message passing model. This is beneficial for modularity, because defining a message-passing format allows us to standardize the physical interconnection between components, a recognized benefit in SoC design [22], and makes it easier to chain offloads together. ClickNP [47] and PANIC [49], two modular packet processing frameworks, have also used a message-based approach. The message passing can be done over dedicated connections, which is the approach used by ClickNP, or a NoC which is used by PANIC.
@@
We prefer a NoC interconnect for two main reasons related to our goal of scalability. First, we can take advantage of the multiplexing provided by the NoC routers. Certain tiles may interact with many other tiles, e.g. if we instantiate multiple copies of the same component or common services such as memory buffer storage. Direct connections can lead to large multiplexers and wires with significant fan-out. Although we could create specialized pipelined multiplexers and arbiters, these essentially look like NoC routers.
@@
Second, we would like the interconnect wiring to remain stable whenever possible. In the ClickNP model, top-level wires are determined by the computational graph. If we wish to form a chain that links together two components that did not communicate before, we must add new interconnect wires, which are typically the longest wires. A NoC allows us to reuse physical wiring to chain any elements that exist in the design, as long as we are careful with deadlock.
These scalability benefits apply both to the data plane and control plane. We discuss the benefits further for the control plane specifically in Section IV-F.
IV-D Tile chain routing